There are 4 traffic lights along the road, each prohibits further movement of the car with a probability of 0.5. Find the

distribution of the number of the traffic lights that car passed to the first stop. What is the variance of the random

variable? 



1
Expert's answer
2022-01-10T14:25:13-0500

Let "X=" the number of the traffic lights that car passed to the first stop.

Then


"P(X=0)=0.5(0.5)^0=0.5"

"P(X=1)=0.5(0.5)^1=0.25"

"P(X=2)=0.5(0.5)^2=0.125"

"P(X=3)=0.5(0.5)^3=0.0625"

"P(X=4)=0.5(0.5)^4=0.03125"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 & 4 \\ \hline
 p & 0.5 & 0.25 & 0.125 & 0.0625 & 0.03125 \\

\end{array}"

"E(X)=0(0.5)+1(0.25)+2(0.125)"

"+3(0.0625)+4(0.03125)=0.8125"

"E(X^2)=0^2(0.5)+1^2(0.25)+2^2(0.125)"

"+3^2(0.0625)+4^2(0.03125)=1.8125"

"Var(X)=E(X^2)-(E(X))^2"

"=1.8125-(0.8125)^2=1.15234375"